OpenStudy (anonymous):

wat is integral of 2 power logx?

OpenStudy (anonymous):

Let u=log x

OpenStudy (anonymous):

2^logx

OpenStudy (anonymous):

logx = e^( log (log(x) )

myininaya (myininaya):

\[\int\limits_{}^{}2^{logx}dx\]

OpenStudy (anonymous):

ya dats true myininaya

myininaya (myininaya):

i was just writing it the way i like to see it lol

OpenStudy (anonymous):

e^(log(2) ) =2

OpenStudy (anonymous):

so 2^log(x) = [ e^(log(2) ) ]^ log(x)

OpenStudy (anonymous):

= e^[(log2)(logx)]

OpenStudy (anonymous):

not sure if its getting anywhere but its similiar to integrating or differentiating something like 2^x

OpenStudy (blacksteel):

You've got the solution right there elec. 2^log(x) = [ e^(log(2) ) ]^ log(x) = [ e^(log(x) ) ]^ log(2) = x^log(2) = x ^(log(2) + 1)/(log(2) + 1)

OpenStudy (blacksteel):

or x*x^log(2) / [1 + log(2)]

OpenStudy (anonymous):

2^logx = x^log2 solve it !!

OpenStudy (anonymous):

\[2^{logx} = x ^{\log2} \] substitute this n ull find the integral of the form \[x ^{?}\]

OpenStudy (anonymous):

blacksteel can u just give me the correct explanation for this??i meant step by step?

OpenStudy (anonymous):

Use this old trick that is usually used for the derivative and apply it to the integral. Let\[y =2^{\log x}\]\[\ln y =(\log x)(\ln 2) \]Take integral both sides. It becomes integral of log x with the coefficient ln 2.